OpenStudy (anonymous):

Metal and fan!!!!!!!!!!!!!! Find the solution of this system of equations. 6x=90+3y x-3y=50

OpenStudy (anonymous):

IT IS SPELLED MEDAL* !!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (campbell_st):

which method do you prefer... elimination, graphical or substitution

OpenStudy (anonymous):

substitution and k

OpenStudy (anonymous):

x=-28 y=-26 6*-28 - 3*-26 = -168 + 78 = -90 now to check the other one

OpenStudy (campbell_st):

so what was your 1st step..?

OpenStudy (anonymous):

Multiply bottom equation by 6, subtract from top, simplify, and get x = 14

OpenStudy (campbell_st):

@MathCerialKiller the aim of the site is to help understanding... and not to give answers..?

OpenStudy (anonymous):

That's ok guys I got it wrong but thanks for trying to help me.

OpenStudy (anonymous):

I'm sorry if that is against the site policy. I just joined the site about a half hour ago. I won't give out answers in the future.

OpenStudy (anonymous):

x=14 was incorrect?

OpenStudy (campbell_st):

FUN = Fantastic Use of Numbers

OpenStudy (anonymous):

Oh, oops: negative 14 not 14

OpenStudy (anonymous):

My negative key doesn't work on my computer;) lol. So i thought I had written negative 14 but it came out as 14.

OpenStudy (anonymous):

Sorry.